### v1

# Analyzing Games in Maker Protocol Part One: A Multi-Agent Influence Diagram Approach Towards Coordination

Abstract
Decentralized Finance (DeFi) ecosystems, exemplified by the Maker Protocol, rely on intricate games to maintain stability and security. Understanding the dynamics of these games is crucial for ensuring the robustness of the system. This motivating research proposes a novel methodology leveraging Multi-Agent Influence Diagrams (MAID), originally proposed by Koller and Milch, to dissect and analyze the games within the Maker stablecoin protocol. By representing users and governance of the Maker protocol as agents and their interactions as edges in a graph, we capture the complex network of influences governing agent behaviors. Furthermore in the upcoming papers, we will show a Nash Equilibrium model to elucidate strategies that promote coordination and enhance economic security within the ecosystem. Through this approach, we aim to motivate the use of this method to introduce a new method of formal verification of game theoretic security in DeFi platforms.
Keywords:
DeFi Multi-Agent Influence Diagrams Maker ProtocolGame Theory Formal Verification Security
1 Introduction
Decentralized Finance (DeFi) [2] has emerged as a transformative force in the financial landscape, offering innovative solutions that challenge traditional centralized systems [1]. At the forefront of this movement lies the Maker Protocol [4], a cornerstone of the DeFi ecosystem, providing the infrastructure for the creation and governance of the Dai stablecoin (see [5]). However, the stability and security of DeFi platforms like Maker heavily rely on intricate coordination games, which necessitate a deep understanding of their dynamics.
In this research, we propose a novel methodology for dissecting and analyzing the games within the Maker Protocol and in turn, similar stablecoin protocols, leveraging Multi-Agent Influence Diagrams (MAIDs) [3]. Originally introduced by Koller and Milch in 2003, MAIDs offer a powerful framework for modeling complex decision-making scenarios involving multiple agents with competing or aligned objectives. By representing users of the Maker protocol as agents and their interactions as edges in a graph, we capture the intricate network of influences that govern user behaviors within the ecosystem.
Our research aims to shed light on the interplay between incentives, behaviors, and outcomes within the Maker Protocol. By employing MAIDs, we aim to systematically analyze the incentives driving user decisions, identify vulnerabilities in the system, and optimize incentive mechanisms to enhance the stability and security of the platform. Furthermore, we will introduce a Nash Equilibrium model [6, 7] in next papers to elucidate strategies that promote coordination among users, fostering economic security and resilience in the DeFi ecosystem.
Through this approach, we not only contribute to a deeper understanding of the Maker Protocol but also provide a methodology that can be extended to other DeFi platforms. By fortifying coordination security and mitigating risks, our research seeks to pave the way for the continued growth and adoption of decentralized financial systems, ushering in a new era of financial inclusivity and innovation. This paper focuses on construction of MAIDs for the Maker protocol and the subsequent iterations would involve employing the ideas of MAID to conduct a full length equilibrium analysis of the protocol using the mathematical ideas presented in the following sections.
2 Background
2.1 Multi Agent Influence Diagrams
Multi Agent Influence Diagrams (MAIDs) combine the ideas of Bayesian Networks (BNs) [8] and Influence Diagrams [9] to describe decision problems related to multiple agents. MAIDs are graphical models used to represent decision-making scenarios involving multiple agents, each with their own objectives, beliefs, and actions. They extend the traditional influence diagram framework to accommodate the complexities of interactions and dependencies among multiple decision-makers. In a MAID, nodes represent variables such as decisions, uncertainties, utilities, and influences, while directed edges indicate causal relationships or dependencies between them. Each agent is associated with a subset of nodes, representing their local decisions, beliefs, and preferences (see [3]). MAIDs allow for the explicit representation of interactions between agents, including cooperation, competition, coordination, and negotiation. This enables the analysis of strategic interactions and the prediction of outcomes resulting from the decisions of multiple agents. There has been a lot of related work done by Google Deepmind (see [10, 11]) on agent incentives for artificial intelligence architectures which forms the underlying motivator to try and bring the same ideas to DeFi.,
2.2 Maker Protocol
The Maker Protocol operates as a decentralized finance (DeFi) platform on the Ethereum blockchain, facilitating the creation of Dai, a stablecoin pegged to the US dollar. As per [4], Users engage with the protocol by depositing cryptocurrency assets, such as Ethereum or other tokens, into smart contracts as collateral. Based on the value of this collateral, users can then generate Dai, typically up to a predetermined percentage of the collateral’s worth. To maintain the stability and security of the system, the protocol enforces various risk parameters, including collateralization ratios and debt ceilings, which dictate the relationship between collateral and the amount of Dai that can be generated. Users pay stability fees on the Dai they generate, serving as interest rates that contribute to the protocol’s stability. Governance of the Maker Protocol is decentralized, with MKR token holders participating in decision-making through a voting mechanism. These holders vote on proposals to adjust risk parameters and other protocol settings, ensuring the protocol remains responsive to market conditions and emerging risks. Additionally, the protocol includes mechanisms for liquidating collateral if its value falls below a certain threshold, helping to mitigate systemic risk and maintain the stability of the Dai ecosystem. Overall, the Maker Protocol offers users a decentralized and trustless means of generating a stable cryptocurrency, providing access to the benefits of DeFi while minimizing counterparty risk.
3 Model
3.1 Mathematical Exposition
As per Koller and Milch [3], to define a Multi-Agent Influence Diagram (MAID), we start with a collection of agents. We represent them as $\textbf{A}=\{A_{i}\}$ for $i\in N$. These agents interact within the Maker Protocol, which is represented by a set of chance variables denoted as $\chi$, and a set of decision variables $D_{A_{i}}$ specific to each agent $A_{i}\in A$.
Chance variables represent uncertain events or outcomes, similar to nodes in Bayesian networks[8], while decision variables are the choices that each agent $A_{i}$ can make, depicted as rectangular nodes in the diagram. We use $D$ to denote $\cup_{A_{i}\in A}D_{A_{i}}$ The agents’ preferences and objectives are captured using utility variables, denoted as $U_{A_{i}}$ for each agent $A_{i}\in A$, where each agent has a set of utility variables. Similar to the decision variables, we use $U$ to denote $\cup_{A_{i}\in A}U_{A_{i}}$. These utility variables quantify the agent’s satisfaction or payoff based on the outcomes of the decision and chance variables. In the Maker Protocol context, chance variables represent factors like market conditions and collateral fluctuations, while decision variables denote actions taken by users and governance entities.
Mathematically, a MAID is represented as a directed acyclic graph (which is used to denote causal relationships, see [12]) where each variable corresponds to a node, and the relationships between variables are depicted through directed edges. For each chance variable $X_{i}\in\chi$, the MAID specifies a conditional probability distribution (CPD) given the instantiations of its parent variables $Pa(N)\subset\chi\cup D$ for each node $N$ in the diagram, which is $P(N|Pa(N))$. Similarly, for each utility variable $U_{A_{i}}$, the MAID specifies a conditional probability distribution, ensuring that the utility value is deterministic based on the values of its parents. The total utility derived by an agent from a particular instantiation of the environment is computed as the sum of the utility values of all relevant utility variables.
Agents determine their actions at decision nodes based on the variables they observe prior to making a decision, typically represented by the parent variables of the decision node. The decision-making process is defined by a set of decision rules, which specify the agent’s strategy or behavior at each decision node. In the context of the Maker Protocol, users and governance entities make decisions based on observed market conditions, collateral fluctuations, and similarly relevant factors to optimize their utility. By utilizing MAIDs, stakeholders can model and analyze the complex interactions and decision-making processes within the Maker Protocol ecosystem to enhance stability, security, and efficiency as we will demonstrate in the next section. The next portion describes the definitions and theorems, from Koller and Milch’s paper [3], which will be useful in our analysis. Readers are encouraged to refer to the original paper to know more about the ideas in detail.
Definition 1
A decision rule for a decision variable $D$ is a function that maps each instantiation $pa$ of parents $\mathcal{P}(D)$ to a probability distribution over the domain $\text{dom}(D)$. An assignment of decision rules to every decision $D_{A_{i}}$ for a particular agent in $A$ is called a strategy.
We define the payoffs at the leaves by computing a distribution over the utility nodes, given an instantiation of $\theta$. For a leaf $\ell$, the payoff for agent $A$ is:
$u_{A}(\ell)=\mathbb{E}_{\theta}[u_{A}\,|\,\text{inst}_{\ell}]$
Theorem 3.1
[…]
Definition 2
A node $D^{\prime}$ is $s$-reachable from a node $D$ in a MAID $K$ if there is some utility node $U_{e}$ such that if a new parent $D^{\prime}$ were added to $D$, there would be an active path in $K$ from $D$ to $U_{e}$ given $\phi(D)\subseteq j\lor\forall r$, where a path is active in a MAID if it is active in the same graph, viewed as a Bayesian Network.
Definition 3
The relevance graph for a MAID $K$ is a directed graph whose nodes are the decision nodes of $K$, and which contains an edge ${}_{\to}D^{\prime}$ if and only if $D^{\prime}$ is $s$-reachable from $D$.
Definition 4
If $K$ is a MAID and $C$ is a strategy profile for $K$, then the joint distribution for $K$ induced by $C$, denoted $Q_{K}^{C}$, is the joint distribution over $\mathcal{X}$ defined by the Bayes network where:
  • The set of variables is $\mathcal{X}$.
  • For each chance variable $X_{i}\in\mathcal{X}$, for each instantiation $\mathbf{pa}$ of parents of $X_{i}$, denoted $Pa(X_{i})$, there is an edge $X_{i}\rightarrow X_{i}$ iff $X_{i}\in\mathbf{pa}$.
  • For each decision variable $D_{i}\in\mathcal{D}$, the CPD for $D_{i}$ is $P(D_{i}|\mathbf{pa})$.
  • For each agent $A_{i}$, the CPD for each decision variable $D_{i}$ is specified by the corresponding decision rule in $C$.
We can now formulate an equation representing the expected utility that agent $A_{i}$ anticipates in a MAID $K$ when the agents adhere to a specific strategy profile $C$. Let $U_{A_{i}}^{C}$ denote the utility, and suppose $U_{A_{i}}^{C}=\sum_{j}u_{A_{i}}^{j}$. Then:
$E[U_{A_{i}}^{C}]=\sum_{\mathbf{v}\in\text{dom}(\mathbf{V})}u_{A_{i}}(\mathbf{v})\prod_{D\in\mathcal{D}}P(D|\mathbf{pa})\quad\text{where}\quad\text{dom}(\mathbf{V})=\text{joint domain of}\ \mathbf{V}.$ (1)
Since the expectation of a sum of random variables equals the sum of the expectations of the individual random variables, we can express this equation as:
$E[U_{A_{i}}^{C}]=\sum_{\mathbf{v}\in\text{dom}(\mathbf{V})}\sum_{j}u_{A_{i}}^{j}P(D|\mathbf{pa})\quad\text{where}\quad\text{dom}(\mathbf{V})=\text{joint domain of}\ \mathbf{V}.$ (2)
After defining the expected utility, we can now elucidate what it signifies for an agent to optimize their decision at one or more of their decision nodes, relative to a given set of decision rules for the other variables.
Definition 5
Let $H$ be a subset of $\mathcal{D}$, and let $C$ be a strategy profile. We define $C_{\text{opt}}$ as optimal for the strategy profile $C$ if, in the induced MAID $K_{(LCoI_{\text{opt}}M)}$, where the only remaining decisions are those in $H$, the strategy $C_{\text{opt}}$ is optimal. Formally, for all strategies $C_{\text{alt}}$:
$EU_{A_{i}}^{C_{\text{opt}}}$ $\geq EU_{A_{i}}^{C_{\text{alt}}}$
$EU_{A_{i}}^{C}$ $\geq EU_{A_{i}}^{C_{\text{opt}}}$
Note that, in this definition, the decision rules assigned to the variables in $H$ by $C$ are not relevant.
In the context of game theory, a strategy profile is deemed to represent rational behavior if it constitutes a Nash equilibrium [7]. Essentially, a strategy profile is a Nash equilibrium if no agent has an incentive to deviate from the strategy prescribed for them by the profile, provided that the other agents adhere to their specified strategies.
Definition 6
A strategy profile $C$ is a Nash equilibrium for a MAID $K$ if, for all agents $A_{i}$, $C_{\text{opt}}$ is optimal for the strategy profile.
3.2 Model Set Up
We analyse the coordination game played between the key stakeholders within the Maker Protocol : the governance DAO and users of the protocol. The dynamics of the game are as follows :
  1. 1.
    Governance entity (DAO), represented as a set of agents $\textbf{G}=\{G_{1},G_{2}\cdots G_{i}\}$, make decisions regarding protocol parameters, such as interest rates, collateral types, and governance policies.
  2. 2.
    Users of the protocol, represented as another set of agent $\textbf{Up}=\{Up_{1},Up_{2}\cdots Up_{i}\}$, make decisions regarding paying back or borrowing more Dai by adding collateral.
Referring [4], in the game between governance and users within the Maker Protocol, users interact with the protocol by borrowing Dai against collateral and governance entities, including voters and decision-makers, shape the protocol’s direction by proposing and voting on changes to parameters and policies. Users are incentivized to maintain adequate collateralization levels and repay debt obligations while governance entities seek to make decisions that enhance protocol stability, security, and efficiency while considering the preferences and welfare of the broader community.
3.3 Multi Agent Influence Diagram Construction
As per [13], the key risk parameters for Maker Vaults are:
  • Debt Ceiling: The Debt Ceiling represents the upper limit on the aggregate debt that can be generated through a singular collateral type within the Maker Protocol. Maker Governance allocates a specific Debt Ceiling to each collateral type, thereby facilitating adequate diversification within the collateral portfolio of the Maker Protocol. Upon reaching its designated Debt Ceiling, further creation of debt becomes unattainable unless existing users repay some or all of their outstanding Vault debt.
  • Stability Fee: The Stability Fee constitutes an annualized interest rate (APY) assessed on the total amount of Dai generated against a Vault’s collateral. This fee is exclusively denominated in Dai.
  • Liquidation Ratio: The Liquidation Ratio signifies the minimum collateral-to-debt ratio required within the Maker Protocol. A lower Liquidation Ratio implies an anticipation of relatively stable collateral prices, while a higher ratio indicates an expectation of greater price volatility. Mathematically, the Liquidation Ratio can be expressed as the ratio of the value of collateral to the total outstanding debt.
  • Liquidation Penalty: The Liquidation Penalty represents an additional fee imposed on the total outstanding Dai generated by a Vault upon liquidation. This penalty serves as an incentive for Vault owners to maintain adequate levels of collateral to mitigate the risk of liquidation. It is typically expressed as a percentage of the total debt outstanding.
  • Collateral Auction Duration: The Collateral Auction Duration specifies the maximum period allocated for the auctioning of collateral associated with Maker Vaults. While debt and surplus auction durations are system-wide parameters, the collateral auction duration is specific to individual Vaults.
  • Auction Bid Duration: The Auction Bid Duration denotes the duration for which an individual bid remains valid before expiring and subsequently closing the auction. This parameter is crucial in determining the efficiency and speed of the auction process within the Maker Protocol.
  • Auction Step Size: The Auction Step Size serves as a risk parameter aimed at incentivizing early participation in auctions and preventing potential abuse by discouraging minimal bid increments. It establishes the minimum bid increment required for successive bids in an auction, promoting a fair and efficient auction mechanism.
3.3.1 Smart contract Modules:
The Maker Protocol relies on various smart contracts to automate its core functionalities. These modules interact with each other and with external actors to maintain system stability and facilitate user interactions. As per [16] :
  • Core Module: The Core Module is crucial to the system as it contains the entire state of the Maker Protocol and controls the central mechanisms of the system while it is in the expected normal state of operation.
  • Collateral Module: The collateral module is deployed for every new ilk (collateral type) added to Vat. It contains all the adapters and auction contracts for one specific collateral type.
  • The System Stabilizer Module: It’s purpose is to correct the system when the value of the collateral backing Dai drops below the liquidation level (determined by governance) when the stability of the system is at risk. The system stabilizer module creates incentives for Auction Keepers (external actors) to step in and drive the system back to a safe state (system balance) by participating in both debt and surplus auctions and, in turn, earn profits by doing so.
  • Oracle Module: An oracle module is deployed for each collateral type, feeding it the price data for a corresponding collateral type to the Vat. The Oracle Module introduces the whitelisting of addresses, which allows them to broadcast price updates off-chain, which are then fed into a median before being pulled into the OSM. The Spot’ter will then proceed to read from the OSM and will act as the liaison between the oracles and dss.
  • MKR Module: The MKR Module contains the MKR token, which is a deployed DS Token contract. It is an ERC20 token that provides a standard ERC20 token interface. It also contains logic for burning and authorized minting of MKR.
  • Governance Module: The Governance Module contains the contracts that facilitate MKR voting, proposal execution, and voting security of the Maker Protocol.The Governance Module has 3 core components consisting of the Chief, Pause and Spell contracts.
  • Rate Module: The Maker Protocol’s Rate Accumulation Mechanism
  • Proxy Module: The Proxy module was created in order to make it more convenient for users/developers to interact with the Maker Protocol. It contains contract interfaces, proxies, and aliases to functions necessary for both DSR and Vault management and Maker governance.
  • Maker Protocol Emergency Shutdown: Shutdown is a process that can be used as a last resort to directly enforce the Target Price to holders of Dai and Vaults, and protect the Maker Protocol against attacks on its infrastructure. Shutdown stops and gracefully settles the Maker Protocol while ensuring that all users, both Dai holders and Vault holders, receive the net value of assets they are entitled to.
The below figure shows the MAID for this game :
Multi Agent Influence Diagram for GovernanceDAO and User game
Deeper Influence Diagram to show interdependencies between parameters on the smart contract level
The MAID model captures the dependencies and influences between the decisions of the governance entity and the users of the Maker Protocol. Nodes in the MAID represent variables such as decisions, uncertainties, utilities, and influences, while directed edges indicate causal relationships or dependencies between them.
The game unfolds over discrete time steps, with each agent making sequential decisions based on their objectives, beliefs, and the actions of the other agent. The decisions of one agent influence the payoffs and strategies available to the other agent, creating a dynamic interplay of incentives and behaviors.
A brief overview of the connections is as follows :
Sustainability
The sustainability of the Maker Protocol hinges on prudent management of risk parameters such as the debt ceiling, liquidation ratio, and stability fee. While higher debt ceilings can potentially increase revenue through stability fees, they also expose the protocol to greater risk if not managed carefully. Similarly, a lower liquidation ratio allows for more borrowing and higher fee generation but escalates the risk of undercollateralization and losses during market downturns. The stability fee, crucial for revenue generation, must strike a delicate balance to avoid discouraging borrowing and reducing demand for DAI. The Governance Module assumes a pivotal role in aligning decisions regarding risk parameters and protocol upgrades with sustainability goals.
Profitability
Profitability within the Maker Protocol ecosystem is intricately linked to the effectiveness of risk parameters and smart contract modules. The stability fee, determined by governance, directly impacts the profitability of Vault owners by influencing the cost of borrowing Dai against collateral. Additionally, the efficacy of the System Stabilizer Module in managing liquidations and surplus auctions affects the profitability of participants, including Keepers engaging in these auctions to earn profits. Higher liquidation penalties incentivize users to maintain adequate collateralization, thereby reducing the need for liquidations and potential losses for the protocol.
Risk Parameters
Risk parameters, governed by the Governance Module, play a pivotal role in capital efficiency, risk mitigation, and profitability within the Maker Protocol ecosystem. Parameters such as the liquidation ratio, debt ceiling, liquidity penalty, and stability fee directly influence user debt, security risk, demand for DAI, and collateral dynamics. These parameters must be meticulously adjusted to ensure high capital efficiency, low risk exposure, and sustainable profitability. For instance, the debt ceiling sets limits on debt generation against specific collateral types to prevent over-leveraging and maintain system stability. Similarly, the liquidation ratio and penalty incentivize users to uphold sufficient collateralization levels, mitigating the risk of liquidation events.
Debt
The debt within the Maker Protocol represents borrowed Dai against provided collateral. Smart contract modules like the Core Module and Rate Module are instrumental in managing user debt by maintaining accurate records and dynamically adjusting stability fees based on market conditions. Risk parameters such as the debt ceiling, liquidation ratio, and penalty further influence debt dynamics by setting limits and incentivizing adequate collateralization.
Demand for DAI
The demand for DAI is intricately linked to borrowing conditions within the Maker Protocol, influenced primarily by the stability fee. Lower stability fees make borrowing DAI more attractive, potentially increasing demand, although this must be balanced with the need for protocol revenue. Additionally, risk parameters such as the liquidation ratio and debt ceiling play crucial roles in managing risk and ensuring the stability of DAI, thus impacting demand. The efficiency of collateral auctions facilitated by modules like the System Stabilizer Module also influences confidence in the protocol and, consequently, demand for DAI.
Security Risks
Security risks within the Maker Protocol stem from vulnerabilities in smart contracts, market manipulation, or unexpected events like oracle failures. The Oracle Module is vital for mitigating security risks by providing accurate price feeds, while the Maker Protocol Emergency Shutdown mechanism serves as a safeguard against critical failures. Adjusting parameters like the liquidation ratio and debt ceiling is essential for managing security risks associated with insolvency or exposure to volatile assets.
Market Conditions
Market conditions significantly impact the stability and performance of the Maker Protocol. Risk parameters such as the liquidation ratio and penalty are designed to account for market fluctuations and maintain system health during adverse conditions. The Oracle Module provides real-time price data, enabling swift reactions to changing market conditions, while dynamically adjusting debt ceilings based on market conditions helps manage risk and maintain stability.
Collateral Dynamics
The Collateral Module manages collateral types and associated parameters within the Maker Protocol. Adjusting risk parameters like the debt ceiling and liquidation ratio influences collateral dynamics by affecting the selection and utilization of collateral assets. Furthermore, the efficiency of collateral auctions, facilitated by modules like the System Stabilizer Module, impacts collateral dynamics by determining the ease of collateral liquidation in case of undercollateralization.
4 Conclusion and Future Work
In this paper, we have successfully quantified the different inter dependencies between the different parameters in the Maker protocol and created a MAID like structure to analyse the games in the protocol. In future iterations, we will analyse the game using the mathematical principles laid down by Koller and Milch and provide a comprehensive view of equilibria in the protocol in different conditions. In subsequent papers, we will also show that MAIDs can be used as a way to assess the incentive security of a defi protocol such as Maker. Using the library developed by Google Deepmind (see [14]), we will develop our own code with a thorough analysis of collateral dynamics (see [15] ), quantify expected values and payoff functions using ideas from Section 3, and we will demonstrate the equilbrium conditions of the protocol in real time so as to motivate a formal verification procedure of game theoretic interactions on Maker protocol.